alexxiona1:

Given the ordered pair P, determine the requested sin, cos, or tan of angle . P(-5, -3) cos =

Nnesha:

94a41d142d424b0393efb530cd742c6f.png

Nnesha:

first draw a dot at given point (-5,-3) and make a right triangle. we know that (x,y)(-5,-3) x=-5 y=-3 we can use Pythagorean theorem to find the length of hypotenuse.

Nnesha:

\[\large\rm a^2+b^2=c^2\] Pythagorean theorem where c is the longest side of right triangle (hypotenuse)

Nnesha:

ba25fbc70bd6460399be0d7e845371ba.png

Nnesha:

\[\rm (-5)^2+(-3)^2=c^2\] solve for c.

Nnesha:

 \[\rm cos(\theta) =\frac{\color{blue}{adjacent}}{\color{green}{hypotenuse}} \] \[\rm sin(\theta) =\frac{\color{red}{opposite}}{\color{green}{hypotenuse}} \] \[\rm tan(\theta) =\frac{\color{red}{opposite}}{\color{blue}{adjacent}} \]

Amythehedgehog:

1 attachment
Amythehedgehog: